OpenStudy (anonymous):

write 60% as afraction in lowest terms

OpenStudy (anonymous):

60 % = 60/100 divide both number by 20 3/5

OpenStudy (anonymous):

60%= 6/10= 3/5 3/5= 60 (in case you need that too)